Solve the following system of equations using the elimination method
2x - 3y =12
-2x + 3y = 12

Answer by Alan3354(69443)

Solve the following system of equations using the elimination method
2x - 3y =12
-2x + 3y = 12
------------------- Add
0 = 24
--> inconsistent
No solution.
----
If you graph them, you get parallel lines.
No intersection, no solution.